OpenStudy (anonymous):

I need help. How do I solve: 4x^2+3=-17

OpenStudy (anonymous):

the square root of -5

OpenStudy (anonymous):

or the square root of 5i make sure the i is not included in the square root symbol thingy XD

OpenStudy (anonymous):

Can you explain how to do it?

OpenStudy (anonymous):

4x^2+3=-17 -3 -3 4x^2 = -20 4x^2 4x^2 x^2 = -5

OpenStudy (anonymous):

Thank you :)